OpenStudy (anonymous):

What is the difference of (x/x-3)-(5x-1/x^2-6x+9) A. x^2-8x-1/(x-3)^2 B. x^2-8x+1/(x-3)^2 C.x^2-2x+1/(x-3)^2 B.x^2-2x-1/(x-3)^2

OpenStudy (luigi0210):

Get a common denominator \[\frac{ x }{ x-3 }-\frac{ 5x-1 }{ (x-3)(x-3) }\]

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

luigi can you pllzz help me with this no one else would and ive been puting in up for an hour straight

OpenStudy (luigi0210):

multiply the first one by x-3

OpenStudy (luigi0210):

\[(\frac{ x }{ x-3 }*\frac{ x-3 }{ x-3 })-\frac{ 5x-1 }{ (x-3)^2 }\]

OpenStudy (luigi0210):

@ineedhelp72

OpenStudy (anonymous):

ok but that isnt one of the answer choices